# ChessGPT: Bridging Policy Learning and Language Modeling

### v1

@@ 1 Introduction
In recent years, large language models (LLMs) based on transformer architectures (vaswani2017attention, ) have showcased remarkable capabilities far exceeding their original design as simple language modeling tools. This was especially notable following the advent of ChatGPT (ouyang2022training, ). Stemming from causal language modeling, a plethora of recent studies have concentrated on developing efficient and powerful LLM base models (du2022glm, ; black2022gpt, ; touvron2023llama, ; biderman2023pythia, ; together2023redpajama, ), supervised fine-tuned models (alpaca, ; vicuna2023, ; gpt4all, ; kopf2023openassistant, ) and models leveraging Reinforcement Learning from Human Feedback (RLHF) (safe_rlhf, ; chatglm, ; moss, ).
@@
Concurrently, there has been a growing trend in employing Large Language Models (LLMs) as foundational elements for decision-making systems. These systems either depend on the expressive capacity of transformer architectures to execute imitation learning, thereby modeling complex behaviors (chen2021decision, ; janner2021offline, ; baker2022video, ), or they harness the common knowledge embedded within LLMs to facilitate the policy learning process (wang2023describe, ; fan2022minedojo, ; du2023guiding, ; ahn2022can, ). However, the dynamic interplay between policy learning and language modeling has been scarcely addressed. Human decision-making typically involves both: we draw upon historical policy interaction to refine our policy and also employ our thoughts for strategic consideration, mostly in natural language form. Based on this logic, we argue that the study of natural language understanding and policy learning should not be isolated. To advance exploration in this realm, we choose one classic game: Chess, as a practical testbed for initial steps in this direction.
@@
Chess, one of the oldest and most universally played board games, presents an ideal testbed due to the wealth of both policy data and language data. In terms of policy data, it is reported that over ten million games are played daily on Chess.com, the most frequented online chess platform. Regarding language data, a myriad of chess-related knowledge is readily accessible online in various forms and mediums, ranging from game analysis, books, puzzles, and news, to online tutorials, wikis, and even YouTube videos. Building on these resources, we have constructed a comprehensive pipeline dedicated to research on chess policy learning and language modeling. Specifically, we provide:
Datasets We curated a large-scale game and language dataset for chess. Our dataset comprises numerous game data from online chess databases recording how humans and chess engines game replay. It also includes a language dataset that encapsulates chess knowledge in a natural language format, as well as a mixed game-language dataset, which offers the most straightforward interrelated data including articles, discussion, or commentary (language) on specific chess game replay (game).
Models We introduce two models, ChessCLIP and ChessGPT, leveraging our datasets. These models showcase the potential for AI to learn from a mixture of replay data and language knowledge.
@@
Evaluations We design an extensive set of tasks to evaluate our models’ abilities from three distinct perspectives: modeling ability, to gauge the model’s proficiency in tracking game state; value judgement ability, measuring the model’s capacity for value assessment and chess knowledge; and policy ability, to test the model’s capability in decision-making. Our experimental results confirm that our models consistently outperform other LLM baselines in all evaluation tasks.
@@
Our work primarily pursues two objectives. Firstly, we construct the whole pipeline on chess as an initial step in promoting research on the interaction/interplay between policy learning and language learning, as well as on the potential of language as a tool for action and understanding. Secondly, our efforts have yielded valuable by-products: the ChessGPT/CLIP models. These models possess practical applicability - they could potentially serve as effective Chess AI assistants for humans.
@@ 2 Related work
There has been increasing interest in leveraging Internet-scale knowledge for creating agents capable of generalizing across many tasks and capabilities (wang2023describe, ; fan2022minedojo, ). For example, MineDojo (fan2022minedojo, ) introduced a framework on Minecraft for understanding how to enable artificial agents to learn in an open-ended environment. We care more about the interplay between language modeling and policy learning.
@@ 3 A large-scale game and language dataset for chess
We introduce a large-scale game & language dataset by collecting all chess-related materials from the Internet. Our dataset can be mainly divided into four categories: (1) The Game dataset, encompassing online chess match replay data involving worldwide human players and chess engines of varying skill levels. (2) The Language dataset, principally recording chess-associated knowledge, analyses, discussions, and news in the form of natural language (3) The Mixed Game-Language dataset, incorporating both game data and human natural language elements (such as game analysis or comments) in alignment. (4) The instruction-tuning and conversation dataset, consisting of instruction data and conversation data related to chess. We include comprehensive descriptions, examples and the procedure of data collection and pre-processing in appendix C.
@@ 3.1 Game dataset
Game replay data provide the most direct method for both humans and machines to grasp the play mechanics of Chess. In chess, these data are commonly stored in the Portable Game Notation (PGN) format which is a standard plain text format as illustrated in fig. 1.
@@
A PGN starts with some headers that include metadata about the game. These headers include information such as the name of players, the Elo ratings, the opening play, and the game outcome. The headers are followed by a move text section that records the moves played by the two players in turn. The moves may be further annotated with comments enclosed in braces.
Lichess dataset We collect five months of online game data from the Lichess database (lichess, ), culminating in a total of 18 million game replay records for online game players.
@@
CCRL Chess engines like StockFish and LeelaZero have attained a proficiency level far beyond what any human player can currently reach. Considering this, we additionally incorporate the Computer Chess Rating Lists (CCRL) (ccrl, ), which is a dataset of chess games played by computer chess engines. The CCRL dataset comprises a considerable collection of chess games, specifically 1.6 million, all of which are played by computer chess engines and stored in PGN format. The Elo-ratings of chess engines fall in the range of 2800-3700.
Chess puzzles A chess puzzle represents a particular chessboard configuration, designed to present a distinct challenge or objective for the solver. Chess puzzles often require players to find the best move or sequence of moves to achieve a specific goal, such as checkmating the opponent’s king, or finding a tactical combination. In our game dataset, we integrate 3.3M puzzles sourced from the Lichess puzzle dataset. Each puzzle within this collection is annotated with its rating, theme and solution.
@@ 3.2 Language dataset
Existing dataset Numerous existing datasets comprise general internet crawl data from platforms like CommonCrawl or Wikipedia. We establish a filtering pipeline to extract only chess-related language corpus from pre-existing language corpus, including C4 (2019t5, ), Pile (gao2020pile, ), Oscar (OrtizSuarezSagotRomary2019, ), Wikipedia (wikidump, ) and RedPajama (together2023redpajama, ). These datasets extend the scope of our language data beyond mere game-play.
Chess blogs Numerous chess websites often publish insightful blogs, sharing their analyses and perspectives on various aspects of chess gameplay. Such blog data is incredibly valuable, as it encompasses game-specific analysis, forming a vital link between the concrete chess game data and its interpretation in natural language form. We manually select approximately 30 chess-related websites and scrape 60k blog articles.
@@
Chess forums Chess forum serves as a platform for a large amount of chess-related dialogues and conversations involving a diverse range of users. These platforms encompass high-quality question-and-answer pairs, as seen in platforms like StackExchange, or more generalized discussions on various chess-related topics, commonly found in dedicated chess-specific forums. We mainly scrape chess forum data from 5 chess-specific forum platforms and StackExchange, using requests and playwright. This process results in a collection of 130K posts, representing a wealth of diverse views, queries, and discourses related to the world of chess.
@@ 3.3 Mixed game-language dataset
These annotated games inherently maintain the correspondence between board state and human language, serving as an exceptionally high-quality data source to align a model with complex human intentions and judgements. We amass annotated games from seven sources, five of which are collected from the internet while the rest two are commercial datasets. In total, we collect 220k annotated games with 1.3M board-language pairs.
Youtube transcripts Drawing inspiration similarly from MineDoJo (fan2022minedojo, ), a YouTube video can naturally serve as a mixed game-language dataset by aligning video clips with natural language transcripts based on timestamps. Rather than generating image-language pairs directly, we develop a pipeline that accurately applies OCR (Optical Character Recognition) to chessboard screenshots to generate FEN (Forsyth-Edwards Notation), a system that describes the chess state in a language format. We gather approximately 40k chess videos, resulting in 100M words in English transcripts and 3M board-language pairs, thus establishing a substantial mixed game-language dataset.
@@ 3.4 Instruction-tuning & conversation dataset
Supervised fine-tuning is a crucial component to train large language model (LLM) to follow instructions (ouyang2022training, ). In addition to the comprehensive chess materials mentioned before, we also collect instruction-tuning and conversation datasets which can be used to finetune the pre-trained LLM base model, thereby enhancing its instruction-following and dialogue capability.
Instruction-tuning data from GPT-4 Inspired by Alpaca (alpaca, ), we use the self-instruct technique (selfinstruct, ) to generate high-quality, instruction-following data through GPT-4 (bubeck2023sparks, ). Specifically, we manually construct 200 seed prompts for chess-related questions or instructions. These prompts serve as few-shot examples, guiding GPT-4 towards more coherent and relevant generation. Finally, we generate around 4k instruction-response pairs using this pipeline.
Conversation data from Reddit The instruction data collected from GPT-4 are mainly in a single-step form, which means only one round of question-answer pair is included. To mitigate this issue, we collect multi-step conversation data about chess on Reddit. Reddit allows users to interact by commenting on posts and responding to other comments, creating a nested structure of responses. This nested structure can be easily converted to a conversation tree by treating the comment’s reply as a child node for that reply. A rich source of conversation data can then be acquired by navigating from the root node to each leaf node via every available path. In all, we choose 6 chess-related sub-reddits and collect 100k human conversations about chess.
@@ 4.1 ChessCLIP
We want to emphasize more on what ChessCLIP can do by aligning the policy modality and the language modality. Firstly, ChessCLIP offers a similarity metric given one PGN and a text description. Just like the application of large-scale image/text retrieval using CLIP, ChessCLIP can help users conduct PGN/text retrieval - search for game based on text or search for comments based on specific game. In addition, because of the low-dimensional feature of action space compared to vision or language space (there only exists a few legal moves for a given chess state ), we can directly conduct search algorithms to maximize the similarity to generate action based on one text description using ChessCLIP. For example, given a chessboard state and a text description, ChessCLIP can generate a move by iterating through all legal moves and finding one move that returns the largest similarity. By the same logic, ChessCLIP can directly generate move sequences (multiple actions) using greedy search or beam search. We refer the reader to section D.1.1 for more discussions.
Implementation details We preprocess the annotated PGNs to produce board/text pairs which we feed separately to the board and text encoders. In particular, for every move in the PGN, we extract the comments attached to the move as well as the board state. We encode the board positions and moves using the same scheme as those used by Leela Chess Zero (lc0) (LeelaChessZero, ), which is similar to the encoding used by AlphaZero (silver2018az, ) for encoding positions and moves in chess. Concretely, the board positions are encoded as a feature map and the actions are encoded as a vector. We instantiate a ChessCLIP model with a pair of text encoder and a board/action encoder. For the text encoder, we only fine-tune the last two layers of pretrained text encoder from OpenAI CLIP model. For the board/action encoder, we use a ResNet (he2016resnet, ) architecture that conditions on the action encoding via a modified FiLM layer (perez2018film, ). Please refer to section D.1.1 for implementation details.
@@ 4.2 ChessGPT
Implementation details We follow common implementations of training a domain-specific instruction-following LLM. Firstly we conduct base-model fine-tuning using chess corpus introduced in section 3.1, 3.2 and 3.3. Due to computational constraints, we choose to finetune the RedPajama-3B-base (together2023redpajama, ) model, which is an open-souce replication of LLaMA (touvron2023llama, ). The base model adopts the GPT-NeoX (black2022gpt, ) architecture, a GPT-3 (brown2020language, ) variant with a few modifications such as rotary positional embedding, parallel attention computation, and different initialization. The base-finetuning brings us our base model: ChessGPT-Base. After base-finetuning, we conduct supervised fine-tuning by supervised learning on question/conversation response using data introduced in section 3.4 and general conversation data from OASST1 (kopf2023openassistant, ), Dolly2 (dolly2, ), Alpaca-GPT4 (peng2023instruction, ), and Sharegpt (sharegpt, ), forming our chat model: ChessGPT-Chat. We leave further RLHF (Reinforcement Learning from Human Feedback) training for future work. Refer to appendix D.1.2 for more details.
@@ 5 Evaluation and benchmark
In this section, we present a comparative analysis between ChessGPT trained on our database with other baseline LLMs. The purpose of our experiments is to assess the performance of ChessGPT in three primary dimensions: Chess modeling ability, Value judgement ability, and Policy ability. The Chess Modeling capability focuses on the language model’s proficiency in accurately tracking the game state and predicting valid moves. Regarding the Value judgement ability, we assess the model’s precision in evaluating the worth of a chess game, encompassing the identification of advantageous positions and the calculation of situation scores. Lastly, the Policy capability gauges the model’s aptitude for generating optimal moves based on a given position. By thoroughly examining these sub-categories, we can comprehensively evaluate and contrast the efficacy of different models in chess-related tasks. We choose the following models as baselines: LLaMA-7B (touvron2023llama, ), RedPajama-Base-3B (together2023redpajama, ), and compare them with ChessCLIP, ChessGPT-Base-3B, and ChessGPT-Chat-3B.

### v2

@@ 1 Introduction
In recent years, large language models (LLMs) based on transformer architectures vaswani2017attention have showcased remarkable capabilities far exceeding their original design as simple language modeling tools. This was especially notable following the advent of ChatGPT ouyang2022training . Stemming from causal language modeling, a plethora of recent studies have concentrated on developing efficient and powerful LLM base models du2022glm ; black2022gpt ; touvron2023llama ; biderman2023pythia ; together2023redpajama , supervised fine-tuned models alpaca ; vicuna2023 ; gpt4all ; kopf2023openassistant and models ouyang2022training ; safe_rlhf ; chatglm ; moss leveraging Reinforcement Learning from Human Feedback (RLHF) (christiano2017deep, ).
@@
Concurrently, there has been a growing trend in employing Large Language Models (LLMs) as foundational elements for decision-making systems. These systems either depend on the expressive capacity of transformer architectures to execute imitation learning, thereby modeling complex behaviors chen2021decision ; janner2021offline ; baker2022video , or they harness the common knowledge embedded within LLMs to facilitate the policy learning process wang2023describe ; fan2022minedojo ; du2023guiding ; ahn2022can . However, the dynamic interplay between policy learning and language modeling has been scarcely addressed. Human decision-making typically involves both: we draw upon historical policy interaction to refine our policy and also employ our thoughts for strategic consideration, mostly in natural language form. Based on this logic, we argue that the study of natural language understanding and policy learning should not be isolated. To advance exploration in this realm, we choose one classic game: Chess, as a practical testbed for initial steps in this direction.
@@
Chess, one of the oldest and most universally played board games, presents an ideal testbed due to the wealth of both policy data and language data. In terms of policy data, it is reported that over ten million games are played daily on Chess.com, the most frequented online chess platform. Regarding language data, a myriad of chess-related knowledge is readily accessible online in various forms and mediums, ranging from game analysis, books, puzzles, and news, to online tutorials, wikis, and even YouTube videos. Building on these resources, we have constructed a comprehensive pipeline dedicated to research on chess policy learning and language modeling, illustrated in fig. 1. Specifically, we provide the following three components:
@@
Datasets. We curated a large-scale game and language dataset for chess. Our dataset comprises numerous game data from online chess databases recording how humans and chess engines game replay. It also includes a language dataset that encapsulates chess knowledge in a natural language format, as well as a mixed game-language dataset, which offers the most straightforward interrelated data including articles, discussion, or commentary (language) on specific chess game replay (game). Models. We introduce two models, ChessCLIP and ChessGPT, leveraging our datasets. These models showcase the potential for AI to learn from a mixture of replay data and language knowledge. Evaluations. We design an extensive set of tasks to evaluate our models’ abilities from three distinct perspectives: modeling ability, to gauge the model’s proficiency in tracking game state; value judgement ability, measuring the model’s capacity for value assessment and chess knowledge; and policy ability, to test the model’s capability in decision-making. Our experimental results confirm that our models consistently outperform other LLM baselines in all evaluation tasks.
@@
We illustrate our full pipeline in fig. 1. Our work primarily pursues two objectives. Firstly, we construct the whole pipeline on chess as an initial step in promoting research on the interaction/interplay between policy learning and language learning, as well as on the potential of language as a tool for action and understanding. Secondly, our efforts have yielded valuable by-products: the ChessGPT/CLIP models. These models possess practical applicability - they could potentially serve as effective Chess AI assistants for humans.
@@ 2 Related work
There has been increasing interest in leveraging Internet-scale knowledge for creating agents capable of generalizing across many tasks and capabilities (wang2023describe, ; fan2022minedojo, ; reed2022generalist, ). For example, MineDojo (fan2022minedojo, ) introduced a framework on Minecraft for understanding how to enable artificial agents to learn in an open-ended environment. More recently, there has been a surge in research that treats LLMs as agents, aiming to harness their internet-scale knowledge for decision-making tasks (liu2023agentbench, ; wang2023voyager, ; yang2023auto, ; hong2023metagpt, ). In contrast to these studies which typically rely on powerful LLMs like GPT-4 (bubeck2023sparks, ), our paper concentrates more on training, especially the interplay between language modeling and policy learning.
@@ 3 A large-scale game and language dataset for chess
We introduce a large-scale game & language dataset by collecting chess-related materials from the Internet. Our dataset can be mainly divided into four categories: (1) The Game dataset, encompassing online chess match replay data involving worldwide human players and chess engines of varying skill levels. (2) The Language dataset, principally recording chess-associated knowledge, analyses, discussions, and news in the form of natural language (3) The Mixed Game-Language dataset, incorporating both game data and human natural language elements (such as game analysis or comments) in alignment. (4) The instruction-tuning and conversation dataset, consisting of instruction data and conversation data related to chess. We include comprehensive dataset descriptions, statistics, and text examples in Appendix D, and the procedure of data collection and pre-processing in Appendix E.
@@ 3.1 Game dataset
Game replay data provide the most direct method for both humans and machines to grasp the play mechanics of Chess. In chess, these data are commonly stored in the Portable Game Notation (PGN111We refer to Appendix B for readers who are not familiar with chess notation format.) format which is a standard plain text format as illustrated in Figure 2. A PGN starts with some headers that include metadata about the game. These headers include information such as the name of players, the Elo ratings, the opening play, and the game outcome. The headers are followed by a move text section that records the moves played by the two players in turn. The moves may be further annotated with comments enclosed in braces.
@@
Lichess dataset We collect five months of online game data from the Lichess database lichess , culminating in a total of 17.5 million game replay records for online game players.
@@
CCRL Chess engines like StockFish and LeelaZero have attained a proficiency level far beyond what any human player can currently reach. Considering this, we additionally incorporate the Computer Chess Rating Lists (CCRL) (ccrl, ), which is a dataset of chess games played by computer chess engines. The CCRL dataset comprises a considerable collection of chess games, specifically 3 million, all of which are played by computer chess engines and stored in PGN format. The Elo-ratings of chess engines fall in the range of 2800-3700.
Chess puzzles A chess puzzle represents a particular chessboard configuration, designed to present a distinct challenge or objective for the solver. Chess puzzles often require players to find the best move or sequence of moves to achieve a specific goal, such as checkmating the opponent’s king, or finding a tactical combination. In our game dataset, we integrate 3.2M puzzles sourced from the Lichess puzzle dataset. Each puzzle within this collection is annotated with its rating, theme and solution.
@@ 3.2 Language dataset
Existing dataset Numerous existing datasets comprise general internet crawl data from platforms like CommonCrawl or Wikipedia. We establish a filtering pipeline to extract only chess-related language corpus from pre-existing language corpus, including C4 (2019t5, ), Pile (gao2020pile, ), Oscar (OrtizSuarezSagotRomary2019, ), Wikipedia wikidump and RedPajama together2023redpajama . These datasets extend the scope of our language data beyond mere game-play.
Chess blogs Numerous chess websites often publish insightful blogs, sharing their analyses and perspectives on various aspects of chess gameplay. Such blog data is incredibly valuable, as it encompasses game-specific analysis, forming a vital link between the concrete chess game data and its interpretation in natural language form. We manually select approximately 30 chess-related websites and scrape 73.2k blog articles.
@@
Chess forums Chess forum serves as a platform for a large amount of chess-related dialogues and conversations involving a diverse range of users. These platforms encompass high-quality question-and-answer pairs, as seen in platforms like StackExchange, or more generalized discussions on various chess-related topics, commonly found in dedicated chess-specific forums. We mainly scrape chess forum data from 5 chess-specific forum platforms and StackExchange, using requests and playwright. This process results in a collection of 140K posts, representing a wealth of diverse views, queries, and discourses related to the world of chess.
@@ 3.3 Mixed game-language dataset
These annotated games inherently maintain the correspondence between board state and human language, serving as an exceptionally high-quality data source to align a model with complex human intentions and judgements. We amass annotated games from seven sources, five of which are collected from the internet while the rest two are commercial datasets. In total, we collect 245K annotated games with 1.3M board-language pairs.
Youtube transcripts Drawing inspiration similarly from MineDoJo (fan2022minedojo, ), a YouTube video can naturally serve as a mixed game-language dataset by aligning video clips with natural language transcripts based on timestamps. Rather than generating image-language pairs directly, we develop a pipeline that accurately applies OCR (Optical Character Recognition) to chessboard screenshots to generate FEN (Forsyth-Edwards Notation), a system that describes the chess state in a language format. We gathered around 83k chess videos, resulting in million-scale English transcripts and board-language pairs, thus establishing a substantial mixed game-language dataset.
@@ 3.4 Instruction-tuning & conversation dataset
Supervised fine-tuning is a crucial component to train large language model (LLM) to follow instructions (ouyang2022training, ; mishra2021cross, ; wei2021finetuned, ). In addition to the comprehensive chess materials mentioned before, we also collect instruction-tuning and conversation datasets which can be used to finetune the pre-trained LLM base model, thereby enhancing its instruction-following and dialogue capability.
Instruction-tuning data from GPT-4 Inspired by Alpaca (alpaca, ), we use the self-instruct technique (selfinstruct, ) to generate high-quality, instruction-following data through GPT-4 bubeck2023sparks . Specifically, we manually construct 200 seed prompts for chess-related questions or instructions. These prompts serve as few-shot examples, guiding GPT-4 towards more coherent and relevant generation. Finally, we generate around 4k instruction-response pairs using this pipeline.
Conversation data from Reddit The instruction data collected from GPT-4 are mainly in a single-step form, which means only one round of question-answer pair is included. To mitigate this issue, we collect multi-step conversation data about chess on Reddit. Reddit allows users to interact by commenting on posts and responding to other comments, creating a nested structure of responses. This nested structure can be easily converted to a conversation tree by treating the comment’s reply as a child node for that reply. A rich source of conversation data can then be acquired by navigating from the root node to each leaf node via every available path. In all, we choose 6 chess-related sub-reddits and collect 410k human conversations about chess.
@@ 4.1 ChessCLIP
We want to emphasize more on what ChessCLIP can do by aligning the policy modality and the language modality. Firstly, ChessCLIP offers a similarity metric given one PGN and a text description. Just like the application of large-scale image/text retrieval using CLIP, ChessCLIP can help users conduct PGN/text retrieval - search for game based on text or search for comments based on specific game. In addition, because of the low-dimensional feature of action space compared to vision or language space (there only exists a few legal moves for a given chess state ), we can directly conduct search algorithms to maximize the similarity to generate action based on one text description using ChessCLIP. For example, given a chessboard state and a text description, ChessCLIP can generate a move by iterating through all legal moves and finding one move that returns the largest similarity. By the same logic, ChessCLIP can directly generate move sequences (multiple actions) using greedy search or beam search. We refer the reader to Section F.1.1 for more discussions.
Implementation details We preprocess the annotated PGNs to produce board/text pairs which we feed separately to the board and text encoders. In particular, for every move in the PGN, we extract the comments attached to the move as well as the board state. We encode the board positions and moves using the same scheme as those used by Leela Chess Zero (lc0) (LeelaChessZero, ), which is similar to the encoding used by AlphaZero (silver2018az, ) for encoding positions and moves in chess. Concretely, the board positions are encoded as a feature map and the actions are encoded as a vector. We instantiate a ChessCLIP model with a pair of text encoder and a board/action encoder. For the text encoder, we only fine-tune the last two layers of pretrained text encoder from OpenAI CLIP model. For the board/action encoder, we use a ResNet (he2016resnet, ) architecture that conditions on the action encoding via a modified FiLM layer (perez2018film, ). Please refer to Section F.1.1 for implementation details.
@@ 4.2 ChessGPT
Implementation details We follow common implementations of training a domain-specific instruction-following LLM. Firstly we conduct base-model fine-tuning using chess corpus introduced in section 3.1, 3.2 and 3.3. Due to computational constraints, we choose to finetune the RedPajama-3B-base together2023redpajama model, which is an open-souce replication of LLaMA touvron2023llama . The base model adopts the GPT-NeoX black2022gpt architecture, a GPT-3 (brown2020language, ) variant with a few modifications such as rotary positional embedding, parallel attention computation, and different initialization. The base-finetuning brings us our base model: ChessGPT-Base. After base-finetuning, we conduct supervised fine-tuning by supervised learning on question/conversation response using data introduced in section 3.4 and general conversation data from OASST1 (kopf2023openassistant, ), Dolly2 (dolly2, ), Alpaca-GPT4 (peng2023instruction, ), and Sharegpt (sharegpt, ), forming our chat model: ChessGPT-Chat. We leave further RLHF (Reinforcement Learning from Human Feedback) training for future work. Refer to Section F.1.2 for more details.
@@ 5 Evaluation and benchmark
In this section, we present a comparative analysis between ChessGPT trained on our database with other baseline LLMs. The purpose of our experiments is to assess the performance of ChessGPT in three primary dimensions: Chess modeling ability, Value judgement ability, and Policy ability. The Chess Modeling capability focuses on the language model’s proficiency in accurately tracking the game state and predicting valid moves. Regarding the Value judgement ability, we assess the model’s precision in evaluating the worth of a chess game, encompassing the identification of advantageous positions and the calculation of situation scores. Lastly, the Policy capability gauges the model’s aptitude for generating optimal moves based on a given position. By thoroughly examining these sub-categories, we can comprehensively evaluate and contrast the efficacy of different models in chess-related tasks. We choose the following models as baselines: LLaMA-7B (touvron2023llama, ), RedPajama-Base-3B (together2023redpajama, ), and compare them with ChessCLIP, ChessGPT-Base-3B222The model size is 2.8B and 3B is a rounded representation, and ChessGPT-Chat-3B. To help readers who are not familiar with chess, we provide task examples and illustrative figures to elucidate these evaluation tasks in Section G.1.